# Critical spin models from holographic disorder

### v1

@@ II.2 Conditions for criticality
Up to this point we have only considered ground states of Gaussian systems, i.e., those that can be mapped to a free-fermion Hamiltonian. This was also the limit of previous matchgate tensor studies and analytical arguments in Ref. Jahn et al. (2022b). This raised the question whether MQA disorder only preserves criticality in the Gaussian setting. To resolve it, we now consider MQA-disordered ground states of the XXZ Hamiltonian (4) in the non-Gaussian regime $\Delta_{0}\neq 0$.
@@
As free-fermion techniques no longer apply, we turn to a numerical tensor network approach to search for the ground state of the Hamiltonian and determine its properties. A widely established method for general 1-dimensional Hamiltonians is given by the Density Matrix Renormalization Group (DMRG) White (1992); Schollwöck (2011). For the following analysis, we apply the DMRG with bond dimension $D=512$ and periodic boundary conditions to approximate the ground state of the disordered XXZ Hamiltonian expressed in a matrix product state (MPS) form. We can calculate the entanglement entropy of relatively small sub-systems ($\ell\leq 8$) by forming the corresponding reduced density matrix out of two copies of the ground state MPS after having traced out the rest of the system. Then we calculate the averaged entanglement entropy as defined in (7).
@@
In Fig. 3, we plot the averaged entanglement entropy for the Gaussian model ($\Delta_{0}=0$) with a black line. This is qualitatively the same result as in Fig.  2(b), but for subsystem size $\ell=6$ and only for $j_{a}/j_{b}<1$. In different shades of orange, we draw the entanglement entropy scaling for the ground states of the non-Gaussian Hamiltonians for different values of the anisotropy constant $\Delta_{0}$. For all values of $\Delta_{0}$ studied, we note that the systems show close-to critical behavior for a range of ratios in the disorder regime. If we consider the case for $\Delta_{0}=1$ (blue line), where now the system is SU$(2)$-symmetric, the entanglement entropy scaling shows similar behavior, with the difference that the width of the criticality plateau is shorter. This difference might be present due to the symmetry transition from U(1) to SU(2) at $\Delta_{0}=1$, which has been a topic of recent interest Lastres et al. (2024); Rylands et al. (2024).
@@
Other quasiperiodic disorder case
@@
$o\rightarrow$ $aaab$ $aaaab$
$a\rightarrow$ $aab$ $aaab$ $ababa$ $aba$ $ab$ $aba$
$b\rightarrow$ $ab$ $aab$ $aba$ $abaaaba$ $a$ $a$
@@
To compare the resulting critical and non-critical phases under various types of disorder, we use a best-fit of $S_{\text{avg}}$ over the whole range of subsystem sizes $\ell$ with respect to the Calabrese-Cardy formula (5) at different system sizes $L$ (an alternative analysis in terms of the behavior of $S_{\text{avg}}$ at fixed $\ell$ is presented in Appendix C). Fig. 4 shows the result of these fits relative to the standard deviation $\sigma$ of the coupling sequences $J_{k}$ of a (non-interacting) XX model Hamiltonian. In addition to the $\{3,7\}$ and $\{4,5\}$ MQA, we also show two examples of non-holographic MQA derived from the Fibonacci and silver-mean inflation rules (shown in Table. 1), as well as two other types of non-MQA quasiperiodic disorder: Coupling sequences from the 2-letter Fibonacci sequence (i.e., the last layer of the corresponding MQA) and the Aubry-André-Harper model Aubry and André (1980); Harper (1955) with couplings
@@
Figure 4: Effective central charge $c$ under disorder, fitted according to the entanglement entropy scaling (5), relative to the standard deviation $\sigma$ of XX model coupling sequences $J_{k}$. The disorder-free $c=1$ phase is denoted by a dashed line. (a)-(b): For the MQA based on the $\{3,7\}$ and $\{4,5\}$ inflations of regular hyperbolic tilings, $c$ remains stable at small $\sigma$ but decays at large $\sigma$. (c)-(d): The Fibonacci and silver-mean MQA, which are not derived from hyperbolic inflation rules, show different behavior: The Fibonacci MQA remains stable at $c\approx 1$ even at large system sizes $L$ and disorder $\sigma$, while in the silver-mean MQA $c$ decays quickly in both limits. (e)-(f): Two examples of non-MQA quasiperiodic sequences: The Fibonacci 2-letter sequence shows decay only at large $L$ and $\sigma$, but is expected to reach an aperiodic singlet phase with $c=0.796$ at $\sigma\to\infty$, while the Aubry-André-Harper model decays quickly towards $c\to 0$ as disorder is increased. For all plots, the curve width represents the parameter error of each fit.
@@
At $\sigma=0$, all of these models in Fig. 4 recover the effective central charge $c=1$ of the translation-invariant critical XX model. Once a disorder with $\sigma>0$ is introduced, however, the three types of sequences (holographic + MQA, non-holographic + MQA, and quasiperiodic non-MQA) show qualitative differences. We find that the $\{3,7\}$ and $\{4,5\}$ MQA both produce a fit with $c\approx 1$ (i.e., equivalent to the translation-invariant case) for a moderate range of disorder even at large $L$, consistent with the results for the $\{3,7\}$ tiling discussed above. And large $\sigma$, both systems transition to a gapped phase with $c\to 0$. In contrast, the silver-mean MQA leads to a $c$ fit that monotonically decays to zero with $\sigma$ even for small system sizes $L$, with the decay rapidly increasing at larger $L$. This indicates the absence of any disordered critical phase in the continuum limit. The Fibonacci MQA, however, does not show a decay of the fitted $c$ at all: At finite $L$, $c$ grows slowly with $\sigma$, while in the $L\to\infty$ limit it approaches a $\sigma$-independent $c=1$. Still, a closer look on the entanglement entropy curve $S_{\text{avg}}(\ell)$ reveals significant deviations from (5) for subsystems of large size $\ell$: As shown in Fig. 5, while for small system sizes the $r,\sigma\to\infty$ limit leads to a $S_{\text{avg}}(\ell)$ scaling apparently following (5) for some values of $c\geq 1$ and $\kappa\approx 1$, at larger inflation steps and system sizes the curve begins to exhibit sections of linear growth with a logarithmic envelope, a characteristic feature of an aperiodic singlet phase described by the SDRG (compare Fig. 2(d)). Such phases appear when most of the entanglement is mediated by (close to) maximally entangled pairs, which indeed appears to be the case here: Spin-spin correlations in the $r\to\infty$ limit become increasingly sparse but remain long-range. This limit strongly deviates from any other MQA we have studied, where the limit of an infinite coupling ratio $r$ always results in a gapped phase; here, we appear to find a transition to another critical phase.
@@
Finally, the two quasiperiodic but non-MQA sequences considered, the Fibonacci 2-letter sequence and the Aubry-André-Harper model, show wildy different behavior. The former leads to an initially increasing $c$ at $\sigma>0$, but converges to a smaller value at large $L$ and $\sigma$. This behavior is well-understood: It corresponds to the strong-disorder limit of an aperiodic singlet phase with an effective central charge Juhász and Zimborás (2007)
@@
Again, this limit leads to an $S_{\text{avg}}(\ell)$ that increases in linear segments, as shown in Fig. 2(d). Unlike the Fibonacci case, Aubry-André-Harper disorder rapidly destabilizes the critical phase for all but very small systems; again, this indicates that there is no stable critical phase at small disorder. Note, however, that a disordered critical point does appear at large disorder $D=2$, where the model undergoes a metal-insulator phase transition Aubry and André (1980); Jitomirskaya (1999); Bu et al. (2022). However, the effective central charge at this critical point differs from the $c=1$ critical point of the XX model Gonçalves (2024).
Figure 5: Averaged entanglement entropy $S_{\text{avg}}$ for the Fibonacci MQA at (a) $n=7$ and (b) $n=11$ inflation steps at different coupling ratios $r=j_{a}/j_{b}$. At large $n$ and $r$, $S_{\text{avg}}$ deviates from a Calabrese-Cardy (CC) fit (5), exhibiting a segmented linear growth typical for an aperiodic singlet phase.
Geometry of non-holographic MQA
@@
Figure 6: (a): The Fibonacci sequence as a mixing of two $\{3,7\}$ aperiodic sequences. (b): A $\{3,7\}$ sequence starting from an initial letter $a$ can be represented on a $\{3,7\}$-hyperbolic lattice as a closed surface, which can cover the whole lattice, following the corresponding inflation steps (blue shaded). On the contrary, $\{3,7\}$ sequence starting from initial letters $ab$ can be represented as a geodesic on a $\{3,7\}$-hyperbolic lattice (orange shaded). (c) The Silver-mean sequence as a mixing of two $\{6,4\}$ aperiodic sequences. (d): Starting with $a$, we can draw a closed surface on a $\{6,4\}$-hyperbolic lattice, which will fill the whole lattice, following the corresponding inflation steps (blue shaded). A sequence starting with $aba$, though, can be represented as a boundary-to-boundary cut on the $\{6,4\}$-hyperbolic lattice (orange shaded).
@@
We now give an interpretation to this behavior in terms of holographic symmetries. While the MQA is a general ansatz class that can be constructed from any inflation rule, those corresponding to the symmetries of regular hyperbolic tilings are endowed with the additional symmetries of the discretized hyperbolic (Poincaré) disk, of the so-called Fuchsian groups Osborne and Stiegemann (2020); Boettcher et al. (2022). This corresponds to a discrete version of the symmetries one finds in holographic dualities such as AdS/CFT, where the $SO(d,2)$ symmetries of $d{+}1$-dimensional anti-de Sitter (AdS) spacetime match the symmetries of a conformal field theory with $d$ spacetime dimensions Brown and Henneaux (1986). Under a regular discretization of AdS time-slices, i.e., the hyperbolic disk, the spatial bulk symmetries that would otherwise match those of CFT ground states become reduced to a Fuchsian group, appearing on the boundary as a set of quasiperiodic symmetries Boyle et al. (2020); Jahn et al. (2020). Theories whose ground states match these more general symmetries (which include the translation-invariant CFT ground states) have been proposed to belong to a new class of quasiperiodic conformal field theories (qCFTs) Jahn et al. (2022a). At small disorder, the $\{p,q\}$ MQA XXZ models considered in this work are candidates for such qCFTs, as their symmetries are expected to match a bulk Fuchsian group and therefore a subset of the symmetries of the $c=1$ critical XXZ model. For the other MQA models, such as the Fibonacci and silver-mean MQA, such symmetries are not guaranteed. As shown in Fig. 6, even and odd layers of both MQAs are associated with different regions on the same $\{p,q\}$ tiling, on which bulk symmetries act differently. Specifically, the letter sequences describing the boundaries of these regions, both of which are periodically repeated, correspond to two different cuts through the hyperbolic disk: Closed curves of finite length centered around a single vertex or tile, leading to a rotational $\mathbb{Z}_{p}$ or $\mathbb{Z}_{q}$ symmetry, or infinitely long cuts between two boundary points with a translational $\mathbb{Z}_{\infty}$ symmetry (see Fig. 7). A bulk transformation preserving one of these symmetries necessarily breaks the other, thus ensuring that the resulting MQA no longer respects qCFT symmetries.
III Discussion
@@
As we have shown, there exists strong numerical evidence that boundary symmetries generated by a bulk hyperbolic geometry – analytically captured by the multiscale quasicrystal ansatz (MQA) – lead to disordered critical phases whose entanglement entropy scaling is consistent with the continuum limit of a conformal field theory (CFT). These critical phases seem to be robust for a range of disorder values, appear in both interacting and non-interacting spin chains, and can be generated from different $\{p,q\}$ hyperbolic tilings. Crucially, the MQA only produces such behavior when it is given a quasiperiodic inflation sequence as an input which corresponds to the symmetries of hyperbolic tiling layers, as we saw in Fig. 4(c, d). As these symmetries have previously appeared in tensor network models of holography, they can properly identified as discrete holographic boundary symmetries. Conversely, non-MQA quasiperiodic disorder, e.g. derived purely from the symmetries of the outermost layer of a hyperbolic tiling, leads to other critical (Fig. 2(d) and Fig. 4(e)) and non-critical (Fig. 4(f)) phases. Hence, boundary theories of discrete-holographic models appear to achieve CFT-like critical behavior only when combining hyperbolic lattice symmetries with an MQA.
While the symmetries of the tensor network construction imply a quasiperiodic conformal field theory (qCFT) in the continuum limit Basteiro et al. (2023), the holographic symmetries studied here are a further restriction leading to a matching between site-averaged correlation functions of the disordered states and those of the non-disordered, translation-invariant one, a behavior previously observed in hyperbolic matchgate tensor networks Jahn et al. (2022b). The precise mathematical relationship between these qCFTs and the CFTs emerging from the non-disordered continuum limit is an interesting question for further study; as such disorder is both apparent on all length scales but preserves average correlators, it seems likely to be describable as a CFT perturbation by a marginal operator in the continuum limit.
@@
Note that even though the MQA was derived from the symmetry of hyperbolic tensor network models, it is still possible to construct such tensor networks to produce non-MQA boundary states Basteiro et al. (2023); however, such constructions then explicitly break the symmetries of the underlying tiling (e.g. by introducing a dependence on the radial direction) such that the qCFT symmetries are no longer upheld. While holographic MQAs appear to exhibit universal critical properties, the non-holographic cases (such as the Fibonacci MQA) lead to other interesting phases that could be studied in greater detail in the future.
@@
The tensor network models considered in this paper do not include bulk degrees of freedom, unlike models of holographic codes Almheiri et al. (2015); Pastawski et al. (2015); Jahn and Eisert (2021). However, such tensor network codes produce boundary (code) states that also exhibit MQA disorder Jahn et al. (2022a), albeit with very sparse correlation functions that do not seem to lead to a smooth continuum theory. However, more sophisticated tensor-network models of AdS/CFT should be capable of combining (approximate) holographic codes that contain (quasiperiodic) CFTs in their code space under a suitable continuum limit. These boundary theories will likely resemble the MQA-disordered models studied here.
@@
Finally, a future challenge will be to implement and study actual physical systems with MQA disorder, for example in quantum simulators with trapped ions Joshi et al. (2023) or Rydberg atoms Yoshida et al. (2024), with the goal of both studying the robustness of MQA criticality under physical noise as well as laying the groundwork for future implementations of discrete-holographic simulations.

### v2

@@ 2.3 Criticality in Non-Gaussian MQA
Up to this point we have only considered ground states of Gaussian systems, i.e., those that can be mapped to a free-fermion Hamiltonian. This was also the limit of previous matchgate tensor studies and analytical arguments in Ref. [19]. This raised the question whether MQA disorder only preserves criticality in the Gaussian setting. To resolve it, we now consider MQA-disordered ground states of the XXZ Hamiltonian (4) in the non-Gaussian regime $\Delta_{0}\neq 0$.
@@
As free-fermion techniques no longer apply, we turn to a numerical tensor network approach to search for the ground state of the Hamiltonian and determine its properties. A widely established method for general 1-dimensional Hamiltonians is given by the Density Matrix Renormalization Group (DMRG) [27, 28]. For the following analysis, we apply the DMRG with bond dimension $D=3000$ and periodic boundary conditions to approximate the ground state of the disordered XXZ Hamiltonian expressed in a matrix product state (MPS) form. We can calculate the entanglement entropy of relatively small sub-systems ($\ell\leq 8$) by forming the corresponding reduced density matrix out of two copies of the ground state MPS after having traced out the rest of the system. Then we calculate the averaged entanglement entropy as defined in (7).
@@
In Fig. 3, we plot the averaged entanglement entropy for the Gaussian model ($\Delta_{0}=0$) with a black line. This is qualitatively the same result as in Fig. 2(c), but for subsystem size $\ell=6$ and only for $j_{b}/j_{a}<1$. In different shades of orange, we draw the entanglement entropy scaling for the ground states of the non-Gaussian Hamiltonians for different values of the anisotropy constant $\Delta_{0}$. For all values of $\Delta_{0}$ studied, we note that the systems show close-to-critical behavior for a range of ratios in the disorder regime. If we consider the case for $\Delta_{0}=1$ (blue line), where now the system is SU$(2)$-symmetric, the entanglement entropy scaling shows similar behavior, with the difference that the width of the criticality plateau is shorter. This difference might be present due to the symmetry transition from U(1) to SU(2) at $\Delta_{0}=1$, which has been a topic of recent interest [29, 30].
Additionally, we calculate the averaged spin-spin correlation decay, defined for the non-Gaussian Hamiltonians ($\Delta_{0}\neq 0$) as:
$c(d)=\frac{1}{L}\sum_{j=1}^{L}\langle\mathbf{S}_{j}\cdot\mathbf{S}_{j+d}\rangle$ (9)
In Fig. 3(b, c) we show the spin-spin correlation decay for two non-Gaussian Hamiltonians, in particular for $\Delta_{0}=0.5$ and $\Delta_{0}=1$, respectively. In both cases, we notice that there exist coupling ratios away from $r=1$ at which we see polynomial correlation decay that is very similar to the critical case at $r=1$. This is in perfect agreement with the averaged entanglement entropy results in Fig. 3(a), discussed above, providing one more piece of evidence that this critical disordered regime is robust even when non-Gaussianity is considered.
@@
Figure 4: Stability of effective central charge $c$ under disorder, fitted according to the entanglement entropy scaling (5), relative to the standard deviation $\sigma$ of XX model coupling sequences $J_{k}$. The disorder-free $c=1$ phase is denoted by a dashed line. (a)-(b): For the MQA based on the $\{3,7\}$ and $\{4,5\}$ regular hyperbolic tilings after $n$ inflation steps with $L$ total sites, $c$ remains stable at small $\sigma$ but decays at large $\sigma$, indicating a breakdown of the critical phase. (c)-(d): The Fibonacci and silver-mean MQA, which are not derived from hyperbolic inflation rules, show different behavior: The Fibonacci MQA remains stable at $c\approx 1$ even at large system sizes $L$ and disorder $\sigma$, while in the silver-mean MQA, $c$ decays quickly in both limits. (e)-(f): Two examples of non-MQA quasiperiodic sequences: The Fibonacci 2-letter sequence shows decay only at large $L$ and $\sigma$, but is expected to reach an aperiodic singlet phase with $c=0.796$ at $\sigma\to\infty$, while the Aubry-André-Harper model decays quickly towards $c\to 0$ as disorder is increased. For all plots, the curve width represents the parameter error of each fit.
2.4 Non-MQA quasiperiodic disorder
@@
[…]
@@
To compare the resulting critical and non-critical phases under various types of disorder, we use a best-fit of $S_{\text{avg}}$ over the whole range of subsystem sizes $\ell$ with respect to the Calabrese-Cardy formula (5) at different system sizes $L$ (an alternative analysis in terms of the behavior of $S_{\text{avg}}$ at fixed $\ell$ is presented in Appendix C). Fig. 4 shows the result of these fits relative to the standard deviation $\sigma$ of the coupling sequences $J_{k}$ of a (non-interacting) XX model Hamiltonian. In addition to the $\{3,7\}$ and $\{4,5\}$ MQA, we also show two examples of non-holographic MQA derived from the Fibonacci and silver-mean inflation rules (shown in Table 1), as well as two other types of non-MQA quasiperiodic disorder: Coupling sequences from the 2-letter Fibonacci sequence (i.e., the last layer of the corresponding MQA) and the Aubry-André-Harper model [31, 32] with couplings
@@
At $\sigma=0$, all of these models in Fig. 4 recover the effective central charge $c=1$ of the translation-invariant critical XX model. Once a disorder with $\sigma>0$ is introduced, however, the three types of sequences (holographic + MQA, non-holographic + MQA, and quasiperiodic non-MQA) show qualitative differences. We find that the $\{3,7\}$ and $\{4,5\}$ MQA both produce a fit with $c\approx 1$ (i.e., equivalent to the translation-invariant case) for a moderate range of disorder even at large $L$, consistent with the results for the $\{3,7\}$ tiling discussed above. And large $\sigma$, both systems transition to a gapped phase with $c\to 0$. In contrast, the silver-mean MQA leads to a $c$ fit that monotonically decays to zero with $\sigma$ even for small system sizes $L$, with the decay rapidly increasing at larger $L$. This indicates the absence of any disordered critical phase in the continuum limit. The Fibonacci MQA, however, does not show a decay of the fitted $c$ at all: At finite $L$, $c$ grows slowly with $\sigma$, while in the $L\to\infty$ limit it approaches a $\sigma$-independent $c=1$. At first glance, this appears to indicate a plateau behavior at small disorder similar to the $\{p,q\}$ MQAs. However, a closer look at the actual entanglement entropy curves $S_{\text{avg}}(\ell)$ shows significant deviations from the translation-invariant case. As shown in Fig. 5(a), it is still possible to fit the data points with a Calabrese-Cardy curve (5) at small system sizes $L$, however only with values of $\kappa$ and $c$ larger than the translation-invariant values (e.g. $\kappa\approx 1.03,c\approx 0.4$ for $r=100$ compared to $\kappa\approx 0.96,c=1/3$ at $r=1$). Fig. 9(c) in the Appendix confirms the absence of an entanglement plateau at small disorder, further distinguishing it from the $\{p,q\}$ MQAs. As we consider the Fibonacci MQA at larger system sizes in Fig. 5(b), we find that $S_{\text{avg}}(\ell)$ starts to deviate from a smooth Calabrese-Cardy curve at large $r$, beginning to exhibit sections of linear growth with a logarithmic envelope, a characteristic feature of an aperiodic singlet phase described by the SDRG (compare Fig. 2(d)). Such phases appear when most of the entanglement is mediated by (close to) maximally entangled pairs, which indeed appears to be the case here: Spin-spin correlations in the $r\to\infty$ limit become increasingly sparse but remain long-range. This limit strongly deviates from any other MQA we have studied, where the limit of an infinite coupling ratio $r$ always results in a gapped phase; here, we appear to find a transition to another critical phase.
@@
Finally, the two quasiperiodic but non-MQA sequences considered, the Fibonacci 2-letter sequence and the Aubry-André-Harper model, show wildy different behavior. The former leads to an initially increasing $c$ at $\sigma>0$, but converges to a smaller value at large $L$ and $\sigma$. This behavior is well-understood: It corresponds to the strong-disorder limit of an aperiodic singlet phase with an effective central charge [21]
@@
Again, this limit leads to an $S_{\text{avg}}(\ell)$ that increases in linear segments, as shown in Fig. 2(d). Unlike the Fibonacci case, Aubry-André-Harper disorder rapidly destabilizes the critical phase for all but very small systems; again, this indicates that there is no stable critical phase at small disorder. Note, however, that a disordered critical point does appear at large disorder $D=2$, where the model undergoes a metal-insulator phase transition [31, 33, 34]. However, the effective central charge at this critical point differs from the $c=1$ critical point of the XX model [35].
Figure 5: Averaged entanglement entropy $S_{\text{avg}}$ for the Fibonacci MQA at (a) $n=7$ and (b) $n=11$ inflation steps at different coupling ratios $r=j_{b}/j_{a}$. Unlike other MQA constructions, entanglement increases monotonically as disorder is increased. At large $n$ and $r$, $S_{\text{avg}}$ also deviates from a Calabrese-Cardy (CC) fit (5), exhibiting a segmented linear growth typical for an aperiodic singlet phase.
Figure 6: (a): The Fibonacci sequence as a mixing of two $\{3,7\}$ aperiodic sequences. (b): A $\{3,7\}$ sequence starting from an initial letter $a$ can be represented on a $\{3,7\}$ hyperbolic lattice as a closed surface, which can cover the whole lattice, following the corresponding inflation steps (blue shaded). On the contrary, a $\{3,7\}$ sequence starting from initial letters $ab$ can be represented as a geodesic on a $\{3,7\}$ hyperbolic lattice (orange shaded). (c) The silver-mean sequence as a mixing of two $\{6,4\}$ aperiodic sequences. (d): Starting with $a$, we can draw a closed surface on a $\{6,4\}$ hyperbolic lattice, which will fill the whole lattice, following the corresponding inflation steps (blue shaded). A sequence starting with $aba$, however, can be represented as a boundary-to-boundary cut on the $\{6,4\}$ hyperbolic lattice (orange shaded).
2.5 Bulk geometry and holography
@@
We now give an interpretation to this behavior in terms of holographic symmetries. While the MQA is a general ansatz class that can be constructed from any inflation rule, those corresponding to the symmetries of regular hyperbolic tilings are endowed with the additional symmetries of the discretized hyperbolic (Poincaré) disk, of the so-called Fuchsian groups [36, 37]. This corresponds to a discrete version of the symmetries one finds in holographic dualities such as AdS/CFT, where the $SO(d,2)$ symmetries of $d{+}1$-dimensional anti-de Sitter (AdS) spacetime match the symmetries of a conformal field theory with $d$ spacetime dimensions [38]. Under a regular discretization of AdS time-slices, i.e., the hyperbolic disk, the spatial bulk symmetries that would otherwise match those of CFT ground states become reduced to a Fuchsian group, appearing on the boundary as a set of quasiperiodic symmetries [10, 11]. Theories whose ground states match these more general symmetries (which include the translation-invariant CFT ground states) have been proposed to belong to a new class of quasiperiodic conformal field theories (qCFTs) [12]. At small disorder, the $\{p,q\}$ MQA XXZ models considered in this work are candidates for such qCFTs, as their symmetries, built from those of a $\{p,q\}$ regular hyperbolic tiling, are expected to match a bulk Fuchsian group. Following the logic above, these symmetries would therefore be a subset of the symmetries of ground states of the critical XXZ model with a $c=1$ continuum CFT limit. For the other MQA models, such as the Fibonacci and silver-mean MQA, such symmetries are not guaranteed. As shown in Fig. 6, even and odd layers of both MQA are associated with different regions on the same $\{p,q\}$ tiling, on which bulk symmetries act differently. Specifically, the letter sequences describing the boundaries of these regions, both of which are periodically repeated, correspond to two different cuts through the hyperbolic disk: Closed curves of finite length centered around a single vertex or tile, leading to a rotational $\mathbb{Z}_{p}$ or $\mathbb{Z}_{q}$ symmetry, or infinitely long cuts between two boundary points with a translational $\mathbb{Z}_{\infty}$ symmetry (see Fig. 7). A bulk transformation preserving one of these symmetries necessarily breaks the other, thus ensuring that the resulting MQA no longer respects qCFT symmetries.
3 Discussion
@@
Understanding the structure of critical boundary theories in holographic bulk-boundary dualities with discretized geometries, such as tensor network models, is an essential step for understanding the relationship between discrete and continuum holography. This discretization necessarily leads to new boundary symmetries not present in the continuum case, suggesting the existence of new disordered critical many-body phases. As we have shown, there exists strong numerical evidence that boundary symmetries generated by a bulk hyperbolic geometry – analytically captured by the multiscale quasicrystal ansatz (MQA) – lead to disordered critical phases whose entanglement entropy scaling is consistent with the continuum limit of a conformal field theory (CFT). These critical phases seem to be robust for a range of disorder values, appear in both interacting and non-interacting spin chains, and can be generated from different $\{p,q\}$ hyperbolic tilings. The appearance of interacting disordered phases is of particular interest here, as they provide a discrete-holographic analogue of boundary theories in continuum holography that are typically considered in the strong coupling regime. Crucially, the MQA only preserves the properties of the translation-invariant model required for a CFT continuum limit — in particular, the same Calabrese-Cardy scaling (5) of entanglement entropy — when it is given a quasiperiodic inflation sequence as an input that also corresponds to the symmetries of hyperbolic tiling layers: As examples for non-hyperbolic MQA that break these symmetries, we saw that a silver-mean MQA does not have stable critical entanglement entropy scaling at increasing disorder (Fig. 4(d)) while a Fibonacci MQA shows immediate deviations from the translation-invariant entanglement entropy scaling with increasing disorder and a transition to a strongly-disordered phase in the infinite-disorder limit. As MQA symmetries derived from hyperbolic tilings have previously appeared in tensor network models of holography, they can be properly identified as discrete-holographic boundary symmetries. Conversely, non-MQA quasiperiodic disorder, e.g. derived purely from the symmetries of the outermost layer of a hyperbolic tiling, leads to other critical (Fig. 2(d) and Fig. 4(e)) and non-critical (Fig. 4(f)) phases. Hence, boundary theories of discrete-holographic models appear to achieve CFT-like critical behavior only when combining hyperbolic lattice symmetries with an MQA. The hyperbolic tensor-network picture appears to suggest that the layers of such an MQA can be understood as a renormalization group transformation beyond the SDRG paradigm. Such an approach and its connection to (discretized) imaginary-time path integrals will be an avenue for future work.
Our numerical results further support the conjecture of Refs. [12, 19] that translation-invariant Hamiltonians with a CFT continuum limit retain their critical properties under a disorder derived from the symmetries of regular hyperbolic tilings, forming a new class of quasiperiodic conformal field theories (qCFT). Intriguingly, we have found here that such behavior even appears without the explicit construction of a tensor network on such hyperbolic tilings, but that boundary theories with the proper (MQA) disorder exhibit qCFT-like critical phases. Studying discrete holography thus appears to lead to distinct many-body phases that are not apparent in models of continuum holography. The precise mathematical relationship between these qCFTs and the CFTs emerging from the non-disordered continuum limit is an interesting question for further study. As the MQA disorder contains contributions on all length scales and does not disappear under coarse-graining, it cannot be described by a CFT perturbation by an irrelevant operator; likewise, our work suggests that average correlations and entanglement of the non-perturbed phase are preserved, ruling out a relevant operator. It thus appears that a qCFT should be describable by a marginal perturbation of a continuum CFT.
@@
Note that even though the MQA was derived from the symmetry of hyperbolic tensor network models, it is still possible to construct such tensor networks to produce non-MQA boundary states [14]; however, such constructions then explicitly break the symmetries of the underlying tiling (e.g. by introducing a dependence on the radial direction) such that the qCFT symmetries are no longer upheld. While holographic MQAs appear to exhibit universal critical properties, the non-holographic cases (such as the Fibonacci MQA) lead to other interesting phases that could be studied in greater detail in the future.
@@
The tensor network models considered in this paper do not include bulk degrees of freedom, unlike models of holographic codes [5, 6, 7]. However, such tensor network codes produce boundary (code) states that also exhibit MQA disorder [12], albeit with very sparse correlation functions that do not seem to lead to a smooth continuum theory. However, more sophisticated tensor-network models of AdS/CFT should be capable of combining (approximate) holographic codes that contain (quasiperiodic) CFTs in their code space under a suitable continuum limit. These boundary theories will likely resemble the MQA-disordered models studied here, as they will obey similar symmetries. General boundary states in these codes will also be non-Gaussian, though some basis states in holographic codes may still be describable as Gaussian states [39].
@@
Finally, a future challenge will be to implement and study actual physical systems with MQA disorder, for example in quantum simulators with trapped ions [40] or Rydberg atoms [41], with the goal of both studying the robustness of MQA criticality under physical noise as well as laying the groundwork for future implementations of discrete-holographic simulations.